# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 5.3. Uniqueness of invariant measures
Appendix A An estimate in Hilbert spaces
The following inequality was noted in [14, Equation (4.9)] without the explicit constant for $\beta>0$.
@@ Appendix A An estimate in Hilbert spaces
In any Hilbert space $H$, we have any $\beta\geq 0$ and for
$C_{\beta}:=((2^{\beta-3}\vee 1)\beta(\beta+2))\vee((2^{\beta-1}\vee 1)(\beta+2)),\quad\beta>0,$
and $C_{0}:=\frac{1}{2}$ that
$0\leq\|x+h\|_{H}^{\beta+2}-\|x\|^{\beta+2}_{H}-(\beta+2)\|x\|^{\beta}_{H}\left\langle x,h\right\rangle\leq C_{\beta}\left(\|x\|_{H}^{\beta}\|h\|_{H}^{2}+\|h\|^{\beta+2}_{H}\right),$
for any $x,h\in H$, where the second inequality is an equality for $\beta=0$ and $C_{0}=\frac{1}{2}$.
Proof.
The first inequality follows from the convex subdifferential inequality, see e.g. [24], noting that the first variation of the functional $F:x\mapsto\|x\|_{H}^{\beta+2}$ is
$\partial F(x)(h)=(\beta+2)\|x\|^{\beta}_{H}\left\langle x,h\right\rangle.$
The second variation of $F$ is given by
$\partial^{2}F(x)(h,h)=\beta(\beta+2)\|x\|^{\beta-2}|\langle x,h\rangle|^{2}+(\beta+2)\|x\|^{\beta}\|h\|^{2}\quad\text{if }x\not=0,$
and $\partial^{2}F(x)(h,h)=0$ if $x=0$.
By Taylor’s formula for Fréchet differentials, see e.g. [79],
$|\|x+h\|_{H}^{\beta+2}-\|x\|^{\beta+2}_{H}-(\beta+2)\|x\|^{\beta}_{H}\left\langle x,h\right\rangle|\leq\frac{1}{2}\partial^{2}F(\xi_{\lambda})(h,h)$
for some $\xi_{\lambda}=\lambda x+(1-\lambda)h$, $\lambda\in(0,1)$. Noting that $\|\xi_{\lambda}\|_{H}\leq\|x\|_{H}+\|h\|_{H}$, we get the claim. The equality for $\beta=0$ follows follows from the polarization identity in Hilbert spaces. ∎
Appendix B A convexity result
(B.1) $g:(x,y)\mapsto|x|^{\alpha}|y|^{\beta}.$
Lemma B.1.
For $\alpha\geq 2$ and $\beta\geq 0$, $g$ as defined in (B.1) is convex.
@@ Appendix B A convexity result
Let $\lambda\in[0,1]$, $x_{1},x_{2},y_{1},y_{2}\in V$. As norms are convex, by Lemma B.1, for $g$ as in (B.1), noting that both components of $g$ are increasing,
@@
Appendix C Small ball probability
We shall need a result on the so-called small ball problem for Lévy processes. As in Hypothesis (G), set
(C.1) $L_{t}:=B\,W_{t}+\int_{D^{c}}G(w)\,\widetilde{N}(t,\mathrm{d}w),\quad t\in[0,T],$
where $(W_{t})_{t\in[0,T]}$ and $\left(\int_{D^{c}}w\,\widetilde{N}(t,\mathrm{d}w)\right)_{t\in[0,T]}$ are independent.
Lemma C.1.
Under Hypotheses (F) and (G), for any $\delta>0$,
(C.2) $\mathbb{P}\left(\sup_{t\in[0,T]}\left\|L_{t}\right\|_{V}<\delta\right)>0,$
where $(L_{t})_{t\in[0,T]}$ is as in (3.4).
Proof.
Fix $T>0$ and $\delta>0$. Let $(L_{t})_{t\in[0,T]}$ have the representation as in Hypothesis (G). For $N\in\mathbb{N}$ denote
$L_{t}^{N}:=\sum_{k=1}^{N}\sqrt{\lambda_{k}}e_{k}l_{t}^{k},$
and let $\lceil T\rceil$ be the smallest integer greater or equal to $T$. By infinite divisibility of $(l^{k})_{s\in[0,T]}$, there exist $\lceil T\rceil$ i.i.d. copies of symmetric Lévy processes $(l^{k,i}_{s})_{s\in\left[0,\frac{T}{\lceil T\rceil}\right]}$, $i=1,\ldots,\lceil T\rceil$ such that
$l_{s}^{k}=\sum_{i=1}^{\lceil T\rceil}l_{\frac{s}{\lceil T\rceil}}^{k,i},\quad s\in[0,T].$
By independence,
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)=\mathbb{P}\left(\sup_{t\in[0,T]}\left\|\sum_{k=1}^{N}\sqrt{\lambda_{k}}e_{k}l_{t}^{k}\right\|_{V}<\frac{\delta}{2}\right)$
$\geq$ $\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=1}^{N}\sqrt{\lambda_{k}}|l_{t}^{k}|\|e_{k}\|_{V}<\frac{\delta}{2}\right)\geq\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=1}^{N}\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|<\frac{\delta}{2}\right)$
$\geq$ $\mathbb{P}\left(\sup_{1\leq k\leq N}\sup_{t\in[0,T]}\left(\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|\right)<\frac{\delta}{2N}\right)\geq\mathbb{P}\left(\bigcap_{k=1}^{N}\left\{\sup_{t\in[0,T]}\left(\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|\right)<\frac{\delta}{2N}\right\}\right)$
$=$ $\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T]}\left(\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|\right)<\frac{\delta}{2N}\right).$
By [7, Corollary 2.1], there exists $K>0$ with
$\begin{split}&-\log\mathbb{P}\left(\sup_{t\in[0,1]}|l_{t}^{k,1}|<\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right)\\ =&K\left[\tilde{\pi}(I^{c}_{k,N,\delta,T})+\frac{4N^{2}\lceil T\rceil^{2}\lambda_{k}\sigma_{k}}{\delta^{2}}\left(\tilde{\sigma}^{2}+\int_{I_{k,N,\delta,T}}u^{2}\,\tilde{\pi}(\mathrm{d}u)\right)\right]=:Ka_{k,N,\delta,T},\end{split}$
where $I_{k,N,\delta,T}:=\left[-\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}},\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right]$, and $(\tilde{\sigma}^{2},\tilde{\pi},0)$ is the characteristic triple of the symmetric Lévy process $(l_{t}^{k,1})_{t\in[0,1]}$, which does not depend on $k$. By Hypothesis (F) we have
$\tilde{\pi}(I_{k,N,\delta,T}^{c})<\infty,\quad\int_{I_{k,N,\delta,T}}u^{2}\,\tilde{\pi}(\mathrm{d}u)<\infty,$
and then,
$\begin{split}&\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T]}|l_{t}^{k}|<\frac{\delta}{2N\sqrt{\lambda_{k}\sigma_{k}}}\right)=\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T]}\left|\sum_{i=1}^{\lceil T\rceil}l_{t/\lceil T\rceil}^{k,i}\right|<\frac{\delta}{2N\sqrt{\lambda_{k}\sigma_{k}}}\right)\\ \geq&\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T/\lceil T\rceil]}|l_{t}^{k,1}|<\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right)\geq\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,1]}|l_{t}^{k,1}|<\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right)\\ =&\exp\left(-K\sum_{k=1}^{N}a_{k,N,\delta,T}\right)\in(0,1),\end{split}$
compare also [43, Lemma 5.6].
By Doob’s submartingale inequality and Hypothesis (G),
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}-L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)=\mathbb{P}\left(\sup_{t\in[0,T]}\left\|\sum_{k=N+1}^{\infty}\sqrt{\lambda_{k}}e_{k}l_{t}^{k}\right\|_{V}^{2}<\frac{\delta^{2}}{4}\right)$
$\geq$ $\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}|l_{t}^{k}|^{2}<\frac{\delta^{2}}{4}\right)=1-\mathbb{P}\left(\left\{\exists t_{0}\in[0,T]\;\colon\;\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}|l_{t_{0}}^{k}|^{2}\geq\frac{\delta^{2}}{4}\right\}\right)$
$\geq$ $1-\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}|l_{t}^{k}|^{2}\geq\frac{\delta^{2}}{4}\right)\geq 1-\frac{4}{\delta^{2}}\mathbb{E}[|l^{1}_{T}|^{2}]\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}.$
Observe that
$01-\frac{4}{\delta^{2}}\mathbb{E}[|l^{1}_{T}|^{2}]\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}>$ 0
if and only if
$\frac{4}{\delta^{2}}\mathbb{E}[|l^{1}_{T}|^{2}]\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}<1.$
Recalling that $\{\lambda_{k}\sigma_{k}\}_{k\in\mathbb{N}}\in\ell^{1}$ by Hypothesis (G), for large enough $N\in\mathbb{N}$,
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}(\omega)-L_{t}^{N}(\omega)\|_{V}<\frac{\delta}{2}\right)>0.$
Hence, by independence,
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}\|_{V}<\delta\right)\geq\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}-L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)\cdot\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)>0.$
@@ Appendix C Small ball probability
Acknowledgments. GB would like to express his gratitude to University of Helsinki, Aalto University School of Science, and Instituto Superior Técnico for all the facilities used along the realization of this work. JMT would like to thank the Department of Mathematics and Systems Analysis, Aalto University for providing its facilities and free coffee for this research. JMT would like to thank Erika Hausenblas (Department of Mathematics, Montanuniversität Leoben, Austria) for providing the important references [7, 43]. JMT would like to thank Mark Veraar (Delft Institute of Applied Mathematics, Delft University of Technology, Netherlands) for an inspiring discussion and for suggesting to apply the stochastic Gronwall inequality from [32]. JMT would like to thank Dirk Blömker (Institut für Mathematik, Universität Augsburg, Germany) for fruitful discussions on stability of the 2D stochastic Navier-Stokes equations.
Funding. The research of GB has been supported by the Finnish Centre of Excellence in Randomness and Structures (decision numbers 346306 and 346308). The research of both authors was partially supported by Horizon Europe Marie Skłodowska-Curie Actions Staff Exchanges (project no. 101183168 – LiBERA).

### v4

@@ 5.3. Uniqueness of invariant measures
6. Mixing times
In this section, we rigorously define the concept of mixing times for uniquely ergodic systems in the Wasserstein distance of order two.
We start recalling the definition of coupling between two probability measures. Let $H$ be a Hilbert space equipped with the inner product $\langle\cdot,\cdot\rangle_{H}$ and its induced norm $\|\cdot\|_{H}$. Let $\mathcal{P}$ be the set of probability measures defined in the measurable space $(H,\mathcal{B}(H))$, where $\mathcal{B}(H)$ is the Borel $\sigma$-algebra of $H$.
Given $\mu_{1}\in\mathcal{P}$ and $\mu_{2}\in\mathcal{P}$, we say that a probability measure $\Pi$ defined in the product space $(H\times H,\mathcal{B}(H)\otimes\mathcal{B}(H))$ is a coupling between $\mu_{1}$ and $\mu_{2}$ if and only if
$\Pi(A\times H)=\mu_{1}(A)\quad\textrm{ and }\quad\Pi(H\times A)=\mu_{2}(A)\quad\textrm{ for any measurable set }\quad A\in\mathcal{B}(H).$
We then denote by $\mathcal{C}(\mu_{1},\mu_{2})$ the set of all couplings between $\mu_{1}$ and $\mu_{2}$. Clearly, $\mathcal{C}(\mu_{1},\mu_{2})\not$$\emptyset$.
Let $\mathcal{P}_{2}$ be the set of Borel probability measures on $(H,\mathcal{B}(H))$ with finite second moment, that is, $\mathcal{P}_{2}\subset\mathcal{P}$ and
(6.1) $\int_{H}\|x\|^{2}_{H}\mu(\mathrm{d}x)<\infty\quad\textrm{ for all }\quad\mu\in\mathcal{P}_{2}.$
The Wasserstein distance of order two, $\mathcal{W}_{2}:\mathcal{P}_{2}\times\mathcal{P}_{2}\to[0,\infty)$, is defined as
(6.2) $\begin{split}\mathcal{W}_{2}(\mu_{1},\mu_{2}):=\left(\inf_{\Pi\in\mathcal{C}(\mu_{1},\mu_{2})}\int_{H\times H}\|x-y\|^{2}_{H}\Pi(\mathrm{d}x,\mathrm{d}y)\right)^{1/2}\end{split}$
for any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}$. By [95, Theorem 6.9] we have that $(\mathcal{P}_{2},\mathcal{W}_{2})$ is a Polish metric space that metrizes the weak topology on $\mathcal{P}_{2}$.
By Theorem 2.1, we have that the random dynamics (2.1) are uniquely ergodic. In other words, there exists a unique probability measure $\mu_{\ast}\in\mathcal{P}$ such that for any initial condition $x\in H$ it satisfies the marginal $X^{x}_{t}$ converges in distribution to $\mu_{\ast}$ as $t\to\infty$. In addition, if $\mu_{\ast}\in\mathcal{P}_{2}$ then $\mathcal{W}_{2}(P^{*}_{t}\delta_{x},\mu_{\ast})\to 0$ as $t\to\infty$. Since the process $(X^{x}_{t})_{t\geq 0}$ is Markovian by [70, Proposition 4.3.5], the map $0,\infty)\ni t\mapsto\mathcal{W}_{2}(P^{*}_{t}\delta_{x},\mu_{\ast})\in[0,\infty)$ is non-increasing, see [12, Lemma B.3 (Monotonicity)].
Let us recall the definition of an $\varepsilon$-mixing time, see for instance [26, 66] in the context of Markov chains. Given a prescribed error $\varepsilon>0$ we define the $\varepsilon$-mixing time (with respect to $\mathcal{W}_{2}$) as
(6.3) $\tau^{x}_{\textsf{mix}}(\varepsilon):=\inf\{t\geq 0:\mathcal{W}_{2}(P^{*}_{t}\delta_{x},\mu_{\ast})\leq\varepsilon\}.$
Lemma 6.1 (Disintegration).
Assume hypotheses and notation of Theorem 2.1. For any $x\in H$ and $0<s\leq t$ it follows that
(6.4) $\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\mu_{\ast})\leq\int_{H}\int_{H}\mathcal{W}_{2}(P^{*}_{s}\delta_{z},P^{*}_{s}\delta_{y})P^{*}_{t-s}\delta_{x}(\mathrm{d}z)\mu_{\ast}(\mathrm{d}y).$
In particular, it follows that
(6.5) $\mathcal{W}_{2}(P^{*}_{t}\delta_{x},\mu_{\ast})=\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\mu_{\ast})\leq\int_{H}\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\delta_{y})\mu_{\ast}(\mathrm{d}y).$
Proof.
Recall Kantorovich’s duality, Theorem 5.10 in [95], for $\mathcal{W}_{2}$, that is,
$\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\mu_{\ast})=\sup_{(f,g)}\int_{H}f(z)P^{*}_{t}\delta_{x}(\mathrm{d}z)-\int_{H}g(z)P^{*}_{t}\mu_{\ast}(\mathrm{d}z),$
where the supremum is running over all continuous and bounded functions $f$ and $g$ such that $f(x)-g(y)\leq\|x-y\|^{2}$ for all $x,y\in H$. Since the family of processes $((X^{x}_{t})_{t\geq 0})_{x\in H}$ is Markovian, for any $0<s\leq t$ we have
(6.6) $\int_{H}f(z)P^{*}_{t}\delta_{x}(\mathrm{d}z)=\int_{H}\mathbb{E}[f(X^{z}_{s})]P^{*}_{t-s}\delta_{x}(\mathrm{d}z)=\int_{H}\int_{H}\mathbb{E}[f(X^{z}_{s})]P^{*}_{t-s}\delta_{x}(\mathrm{d}z)P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)$
and
(6.7) $\int_{H}g(z)P^{*}_{t}\mu_{\ast}(\mathrm{d}z)=\int_{H}\mathbb{E}[g(X^{y}_{s})]P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)=\int_{H}\int_{H}\mathbb{E}[g(X^{y}_{s})]P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)P^{*}_{t-s}\delta_{x}(\mathrm{d}z).$
We then observe that
(6.8) $\begin{split}&\int_{H}f(z)P^{*}_{t}\delta_{x}(\mathrm{d}z)-\int_{H}g(z)P^{*}_{t}\mu_{\ast}(\mathrm{d}z)\\ &=\int_{H}\int_{H}\mathbb{E}[f(X^{z}_{s})]P^{*}_{t-s}\delta_{x}(\mathrm{d}z)P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)-\int_{H}\int_{H}\mathbb{E}[g(X^{y}_{s})]P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)P^{*}_{t-s}\delta_{x}(\mathrm{d}z)\\ &=\int_{H}\int_{H}\left(\mathbb{E}[f(X^{z}_{s})]-\mathbb{E}[g(X^{y}_{s})]\right)P^{*}_{t-s}\delta_{x}(\mathrm{d}z)P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)\\ &\leq\int_{H}\int_{H}\left(\mathbb{E}[f(X^{z}_{s})]-\mathbb{E}[g(X^{y}_{s})]\right)P^{*}_{t-s}\delta_{x}(\mathrm{d}z)P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y).\end{split}$
Now, taking the supremum is running over all continuous and bounded functions $f$ and $g$ such that $f(x)-g(y)\leq\|x-y\|^{2}$ for all $x,y\in H$ in both sides of (6.8) we obtain
(6.9) $\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\mu_{\ast})\leq\int_{H}\int_{H}\mathcal{W}_{2}(P^{*}_{s}\delta_{z},P^{*}_{s}\delta_{y})P^{*}_{t-s}\delta_{x}(\mathrm{d}z)P^{*}_{t-s}\mu_{\ast}(\mathrm{d}y)$
for all $0<s\leq t$, which concludes the proof of (6.4). Since $P^{*}_{t-s}\mu_{\ast}=\mu_{\ast}$ for all $0\leq s\leq t$ and $P^{*}_{0}\delta_{x}=\delta_{x}$, we deduce
(6.10) $\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\mu_{\ast})\leq\int_{H}\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\delta_{y})\mu_{\ast}(\mathrm{d}y)$
yielding the proof of (6.5). ∎
6.1. Proof of Theorem 2.2
The following theorem provides a quantitative upper bound for the $\varepsilon$-mixing time.
Theorem 6.2 (Mixing time).
Assume Hypotheses (A), (B), (C), (D) and (F), (G), (H), and (I) hold true. Suppose that $\eta\equiv 0$ or $\rho\equiv 0$. Let $\gamma\in(0,\delta_{2}]$ be as in Hypothesis (I). Then for any $x\in H$ and $\varepsilon>0$ it follows that the $\varepsilon$-mixing time $\tau^{x}(\varepsilon)$ satisfies the following non-asymptotic estimate:
(6.11) $\tau^{x}_{\textup{{mix}}}(\varepsilon)\leq\frac{2}{\gamma}\left[\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\|x\|^{\beta+2}_{H}+\log\left(\|x\|_{H}+\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)^{1/(\alpha+\beta)}\right)+\log\left(\frac{1}{\varepsilon}\right)\right].$
Proof.
Note that by Hypothesis (I), by Lemma 6.1, and similar ideas as in the proof of Proposition 5.3 and Proposition 5.1, and (5.24) in Lemma 5.9, we have
(6.12) $\begin{split}\mathcal{W}_{2}(P^{*}_{t}\delta_{x},\mu_{\ast})\\ \leq&\int_{H}\mathcal{W}_{2}(P^{*}_{t}\delta_{x},P^{*}_{t}\delta_{z})\mu_{\ast}(\mathrm{d}z)\\ \leq&\int_{H}\left(\mathbb{E}\|X_{t}^{x}-X_{t}^{z}\|_{H}^{2}\right)^{1/2}\mu_{\ast}(\mathrm{d}z)\\ \leq&\exp\left(-\frac{\delta_{2}}{2}t+\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\left(\|x\|^{\beta+2}_{H}+(c_{1}+c_{2}+c_{3})t\right)\right)\int_{H}\|z-x\|_{H}\mu_{\ast}(\mathrm{d}z)\\ \leq&\exp\left(-\frac{\gamma}{2}t+\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\|x\|^{\beta+2}_{H}\right)\int_{H}\|z-x\|_{H}\mu_{\ast}(\mathrm{d}z)\\ \leq&\exp\left(-\frac{\gamma}{2}t+\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\|x\|^{\beta+2}_{H}\right)\|x\|_{H}\\ &+\exp\left(-\frac{\gamma}{2}t+\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\|x\|^{\beta+2}_{H}\right)\int_{H}\|z\|_{H}\mu_{\ast}(\mathrm{d}z)\\ \leq&\exp\left(-\frac{\gamma}{2}t+\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\|x\|^{\beta+2}_{H}\right)\left(\|x\|_{H}+\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)^{1/(\alpha+\beta)}\right)\\ &=:h(t),\end{split}$
for any $x\in H$ and $t\geq 0$. Note that function $h$ is strictly decreasing. Let $\varepsilon>0$ be fixed and let $t_{\varepsilon}>0$ be such that $h(t_{\varepsilon})=\varepsilon$, that is,
(6.13) $t_{\varepsilon}:=\frac{2}{\gamma}\left[\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\|x\|^{\beta+2}_{H}+\log\left(\|x\|_{H}+\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)^{1/(\alpha+\beta)}\right)+\log\left(\frac{1}{\varepsilon}\right)\right]$
By the definition of mixing time (6.3) we obtain $\tau^{x}_{\textsf{mix}}(\varepsilon)\leq t_{\varepsilon}$ and hence (6.11). The proof is complete. ∎
Thus we have proved Theorem 2.2. As a result we obtain the exponential ergodicity of the semigroup $(P_{t})_{t\geq 0}$.
Appendix A A convexity result
(A.1) $g:(x,y)\mapsto|x|^{\alpha}|y|^{\beta}.$
@@ Appendix A A convexity result
For $\alpha\geq 2$ and $\beta\geq 0$, $g$ as defined in (A.1) is convex.
@@
Let $\lambda\in[0,1]$, $x_{1},x_{2},y_{1},y_{2}\in V$. As norms are convex, by Lemma A.1, for $g$ as in (A.1), noting that both components of $g$ are increasing,
@@
Acknowledgments. GB would like to express his gratitude to University of Helsinki, Aalto University School of Science, and Instituto Superior Técnico for all the facilities used along the realization of this work. JMT would like to thank the Department of Mathematics and Systems Analysis, Aalto University for providing its facilities and free coffee for this research. JMT would like to thank Dirk Blömker (Universität Augsburg, Germany) and Martin Hairer (Imperial College, UK and EPFL, Switzerland) for fruitful discussions and useful comments on stability of the 2D stochastic Navier-Stokes equations. JMT would like to thank Erika Hausenblas (Montanuniversität Leoben, Austria) and Mark Veraar (Delft University of Technology, Netherlands) for comments on a previous version of this work.
Funding. The research of GB has been supported by the Finnish Centre of Excellence in Randomness and Structures (decision numbers 346306 and 346308). The research of both authors was partially supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Actions Staff Exchanges (Grant agreement no. 101183168 – LiBERA, Call: HORIZON-MSCA-2023-SE-01).
Disclaimer. Funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Education and Culture Executive Agency (EACEA). Neither the European Union nor EACEA can be held responsible for them.